A flat circular coil with 40 loop of wire has a diameter of 32cm what current most flow in it's wire to produce a field in air of 3.0x10^-4 w/m2 or centre

1
Expert's answer
2022-02-01T09:57:42-0500

Answer

No. Of loops N=40

Radius R=0.16m

Magnetic field on centre

"B=3.0*10^{-4 }W\/m^2"

The current flow in loop is

"I=\frac{RB4\pi}{N\mu_0 2\pi }"

Putting all values

"I=\frac{0.16\times3\times10^{-4}}{40\times10^{-7}\times2\times3.14}"

"I=1.91A."

So current in wire is 1.91A.